# CTD4 - A Deep Continuous Distributional Actor-Critic Agent with a Kalman Fusion of Multiple Critics

### v2

@@ 5 Experiments
We purposefully avoid applying or comparing our method with previous approaches, such as C51 or IQN, which are tailored for discrete action spaces. Our research squarely focuses on addressing the challenges posed by complex tasks characterized by continuous action spaces—a domain notorious for its heightened difficulty level.
To conduct experiments with our proposed method, we select the challenging continuous control tasks from the DeepMind Control Suite (DMCS) (Tassa et al., 2018). DMCS is an established and widely recognized RL test-bed, offering complex environments custom-made for continuous control tasks. We choose ten environments, each presenting distinct levels of complexity and task requirements. See Figure 7 for a graphic representation of each environment.
@@
The default reward signal provided for each environment ensures consistency and comparability across evaluations. The experiments are conducted over a training duration of $1\times 10^{6}$ steps for each environment, employing identical hyperparameters for five independently seeded runs. We regularly evaluate the agent’s performance in each environment during the training; after every $1\times 10^{4}$ training steps, we compute the average reward over $10$ consecutive episodes. It is important to note that no gradient updates are executed during the evaluation phase. Notably, all model parameters are optimized using the Adam optimizer, with a single gradient update executed per environment step. The training process was completed on a PC with an i9 CPU, 128GB of RAM, and a GeForce RTX4090 NVIDIA GPU with Ubuntu 20.04. For a comprehensive list of our chosen hyperparameters, please refer to Appendix A.
@@
To facilitate the replication of our work, we provide all raw data plots corresponding to each independent seed, a comprehensive list of hyperparameters, the entire source code, and task demonstration videos. This supplementary material is available on the paper’s website at: https://sites.google.com/view/ctd4
6 Results and Discussion
@@
Figure 6: Learning curves for CTD4, REDQ and TD3 on ten DeepMind Control Suite continuous control tasks. In each plot, the solid line represents the average, while the shaded region represents the standard deviation of the average evaluation over five independent seeds. Curves have not been smoothed to demonstrate clarity in our results
Following thorough testing and analysis, the results presented in Figure 6 illustrate the average evaluation reward across multiple seeds. Each plot serves to visually convey and compare the performance of our proposed CTD4 against TD3 and REDQ under identical conditions.
@@
However, in the spirit of transparency regarding the results, it is important to acknowledge that in a specific environment, “Cheetah Run,” the proposed method employing continuous distributions fails to surpass TD3. While the performance is reasonably similar, TD3 exhibits a marginal superiority in this particular case. A similar scenario occurs with the “Finger Turn Hard” environment, where REDQ narrowly surpasses CTD4. However, TD3 shows the worst performance in this setting.
This discrepancy may arise due to distinctive behaviours exhibited by Cheetah Run and Finger Turn Hard compared to other environments in the test bed. Subtleties in the reward distribution relative to certain actions might not have been explicitly considered in the approach. Further investigation is warranted to understand the specific dynamics at play in these two environments, which might require tailored adjustments to enhance the performance of the proposed method in these particular settings.
@@ Conclusions
Categorical Distributional Reinforcement Learning has demonstrated efficacy and enhanced the performance of traditional RL. Nonetheless, its practical implementation is complex, marked by numerous computations and tuning steps. A noteworthy limitation in the existing literature is the inability of several categorical distributional RL approaches to address complex tasks within continuous action spaces effectively. In this paper, we set out to investigate the use of continuous distributions as a mechanism for improving and simplifying the use of distribution in RL. Our proposed solution involves leveraging continuous distributions, specifically a normal distribution parameterized by mean and standard deviation. This approach addresses and mitigates the majority of the problems associated with categorical distributional RL, facilitating a more straightforward implementation process. Importantly, our method bypasses the need for custom loss functions, as we minimize the Kullback-Leibler divergence between two distributions, a reasonably manageable loss metric to minimize.
@@
Appendix A Hyperparameters
Table 1: Complete list of Hyperparameters employed in our proposal. TD3 and REDQ values are based on the official implementation provided by the author.
TD3(Fujimoto et al., 2018) Ours REDQ(Chen et al., 2021)
Hyperparameter Value Value Value
Actor Learning Rate $3.10^{-4}$ $10^{-4}$ $3.10^{-4}$
Critic Learning Rate $3.10^{-4}$ $10^{-3}$ $3.10^{-4}$
Optimizer ADAM
Discount Factor $\gamma$ 0.99
Number of Critics $N$ 2 3 10
Batch Size 256
Target Update Rate $\tau$ $5.10^{-3}$
Targets and Actor Update Interval 2
Gradients per Steps $G$ 1
Replay Buffer Size $D$ $10^{6}$
Exploration Policy Noise $\varepsilon$ $\mathcal{N}(0,0.1)$
Decrease Noise per Step
Decrease rate $9.10^{-5}$
Seed Tested 10, 25, 35, 45, 55
Appendix B Training algorithm
Algorithm 1 CTD4
  Initialize: critics networks $Q_{\theta n}$ for $n\in[1..N]$ and actor network $\pi_{\phi}$
  Initialize: target critics networks $\theta^{\prime}n$leftarrow $\theta n$ for $n\in[1..N]$ and target actor network $\phi^{\prime}$leftarrow $\phi$
  Initialize: replay memory $D$
  for $t$ to $T$ do
     Select action $a_{t}\sim\pi_{\phi}(s)+$ decreasing exploration noise $\varepsilon\sim\mathcal{N}(0,\sigma)$
     Store transition tuple $(s_{t},a_{t},r_{t},s_{t+1})$ in $D$
     for $G$ gradient step do
        Sample a mini-batch of $M$ transitions from $D$
[…]
        $\mu_{n}^{\prime},\sigma_{n}^{\prime}\leftarrow Q_{\theta^{\prime}n}(s_{t+1},\tilde{a})$ for $n\in[1..N]$
        $\mu_{k},\sigma_{k}\leftarrow$ Kalman $(\mu_{n}^{\prime},\sigma_{n}^{\prime})$
        $y\leftarrow\mathcal{N}(r+\gamma\mu_{k},\gamma\sigma_{k})$
        $\mu_{n},\sigma_{n}\leftarrow Q_{\theta n}(s_{t},a)$ for $n\in[1..N]$
        Update critics with TD error
        $\theta_{n}\leftarrow argmin_{\theta n}$ KL divergence $(y-\mathcal{N}(\mu_{n},\sigma_{n}))$
        if $t$ mod $d$ then
           $a_{act}\leftarrow\pi_{\phi}(s_{t})$
           $\mu_{n},\sigma_{n}\leftarrow Q_{\theta n}(s_{t},a_{act})$ for $n\in[1..N]$
           $\mu_{k},\sigma_{k}\leftarrow$ Kalman $(\mu_{n},\sigma_{n})$
           Update actor with Deterministic Policy Gradient
           $\nabla_{\phi}J(\phi)\approx\frac{1}{M}\sum_{i=1}^{M}\nabla_{a}\ \mu_{k}\nabla_{\phi}\pi_{\phi}(s_{t})$
           Soft update target networks
           $\theta^{\prime}_{n}\leftarrow\tau\theta_{n}+(1-\tau)\theta^{\prime}_{n}$ for $n\in[1..N]$
           $\phi^{\prime}_{n}\leftarrow\tau\phi_{n}+(1-\tau)\phi^{\prime}_{n}$ for $n\in[1..N]$
        end if
     end for
  end for
Appendix C DMCS Continuous Control Tasks
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 7: The ten continuous control tasks from the DeepMind Control Suite used in our experimentation. From top to bottom and from left to right. Acrobot Swingup, Ball in Cup Catch, Cartpole Swingup, Cheetah Run, Finger Turn Hard, Fish Swim, Hooper Hop, Humanoid Run, Reacher Hard, Walker Walk

### v3

@@ Introduction
We purposefully avoid applying or comparing our method with previous approaches, such as C51 or IQN, which are tailored for discrete action spaces. Our research squarely focuses on addressing the challenges posed by complex tasks characterized by continuous action spaces—a domain notorious for its heightened difficulty level. Additionally, we do not compare our proposal against TD4 and DSAC, as they rely on parallelized learning. Instead, our focus is on improving learning efficiency in a single-agent setup, which is more practical and resource-efficient.
@@
To conduct experiments with our proposed method, we select the challenging continuous control tasks from the DeepMind Control Suite (DMCS) (Tassa et al. 2018). We choose ten environments, each presenting distinct levels of complexity and task requirements. The default reward signal provided for each environment ensures consistency and comparability across evaluations. The experiments are conducted over a training duration of $1\times 10^{6}$ steps for each environment, employing identical hyperparameters for five independently seeded runs. We regularly evaluate the agent’s performance in each environment during the training; after every $1\times 10^{4}$ training steps, we compute the average reward over $10$ consecutive episodes. It is important to note that no gradient updates are executed during the evaluation phase. Notably, all model parameters are optimized using the Adam optimizer, with a single gradient update executed per environment step. The training process was completed on a PC with an i9 CPU, 128GB of RAM, and a GeForce RTX4090 NVIDIA GPU with Ubuntu 20.04.
@@
To facilitate the replication of our work, we provide all raw data plots corresponding to each independent seed, a comprehensive list of hyperparameters, a training loop, the entire source code, and task demonstration videos. This supplementary material is available on the paper’s website at: https://sites.google.com/view/ctd4.
Results and Discussion
@@
Figure 6: Learning curves for CTD4, REDQ and TD3 on ten DeepMind Control Suite continuous control tasks. In each plot, the solid line represents the average, while the shaded region represents the standard deviation of the average evaluation over five independent seeds.
Figure 6 shows the average evaluation reward across multiple seeds, comparing the performance of CTD4, TD3, and REDQ under identical conditions.
@@ Conclusions
CDRL has demonstrated efficacy and enhanced the performance of traditional RL. Nonetheless, its practical implementation is complex, marked by numerous computations and tuning steps. A noteworthy limitation in the existing literature is the inability of several CDRL approaches to address complex tasks within continuous action spaces effectively. In this paper, we set out to investigate the use of continuous distributions as a mechanism for improving and simplifying the use of distribution in RL. Our proposed solution involves leveraging continuous distributions, specifically a normal distribution parameterized by mean and standard deviation. This approach addresses and mitigates the majority of the problems associated with CDRL, facilitating a more straightforward implementation process. Importantly, our method bypasses the need for custom loss functions, as we minimize the KL divergence between two distributions, a reasonably manageable loss metric to minimize.
@@
Acknowledgements
This research was partially supported by the New Zealand Ministry for Business, Innovation and Employment (MBIE) on contract UOAX1810.